# Beyond the random phase approximation for calculating Curie temperatures in ferromagnets: application to Fe, Ni, Co and monolayer CrI3

### v1

@@ A.1 Holstein-Primakoff bosonization
For the HP method, one introduces a bosonic transformation for each site in the unit cell.
@@
$J^{ab}_{\mathbf{q}}$ $=\sum_{i}J^{ab}_{0i}e^{i\mathbf{q}\cdot\mathbf{R}_{i}}$ (56)
@@ A.2 Random Phase Approximation
With multiple sites in the unit cell we consider the retarded Green function
@@
where the commutator yields:
@@
The Green function is thus be obtained as
@@ A.3 Callen Decoupling
For the Callen decoupling scheme the higher order Green function is approximated by
@@
$+\frac{\langle S^{z}\rangle}{2S^{2}N_{q}}\sum_{\mathbf{q^{\prime}}}\Bigl{\{}\delta_{ab}\sum_{c}J^{ac}_{\mathbf{q^{\prime}}}\psi^{ca}_{\mathbf{q^{\prime}}}-(\lambda^{ab}_{\mathbf{q-q^{\prime}}}+J^{ab}_{\mathbf{q-q^{\prime}}})\psi^{ab}_{\mathbf{q^{\prime}}}\Bigr{\}}$
@@
The Green’s function has the same structure as Eq. (68) but with eigenvalues and eigenvectors obtained by diagonalizing Eq. (A.3).

### v2

@@ A.1 Holstein-Primakoff bosonization
For the HP method, one introduces a bosonic transformation for each site in the unit cell:
@@
$J^{ab}_{\mathbf{q}}$ $=\sum_{i}J^{ab}_{0i}e^{i\mathbf{q}\cdot\mathbf{R}_{i}},$ (56)
@@ A.2 Random Phase Approximation
With multiple sites in the unit cell we consider the retarded Green’s function
@@
where the commutator yields
@@
The Green’s function is thus be obtained as
@@ A.3 Callen Decoupling
For the Callen decoupling scheme the higher order Green’s function is approximated by
@@
$+\frac{\langle S^{z}\rangle}{2S^{2}N_{q}}\sum_{\mathbf{q^{\prime}}}\Bigl{\{}\delta_{ab}\sum_{c}J^{ac}_{\mathbf{q^{\prime}}}\psi^{ca}_{\mathbf{q^{\prime}}}-(\lambda^{ab}_{\mathbf{q-q^{\prime}}}+J^{ab}_{\mathbf{q-q^{\prime}}})\psi^{ac}_{\mathbf{q^{\prime}}}\Bigr{\}}$
@@
The Green’s function is again given by Eq. (68) but with $\omega_{\mathbf{q}n}$ being the eigenvalues of Eq. (A.3), which are diagonalized by the corresponding $U_{\mathbf{q}an}$.